OpenStudy (anonymous):

Solve the formula P = 2l + 2w for w.

OpenStudy (anonymous):

are you asking what is w

OpenStudy (anonymous):

P = 2L + 2W Subtract 2L from both sides: P - 2L = 2L - 2L + 2W 2L - 2L is 0, so this simplifies P - 2L = 2W Now divide both sides by 2 (P - 2L)/2 = 2W/2 2/2 equals 1, which you can leave out in multiplication: (P - 2L)/2 = W That's your answer. W = (P - 2L)/2

OpenStudy (anonymous):

Yes .. w=?

OpenStudy (anonymous):

P = 2W +2L step is keep 2L in one side of equation , and move every thing else to the other side. In moving any quantity from one side to other, change the sign. ( equivalent of adding or subtracting equal quantity from both sides of equation). 2L = P - 2W divide both sides by 2: L = p/2 - w Step is the same as solving X in equation of 8 = 2X + 16 2X = 8 -16 = 8 X =4 . But, here you solve one variable in terms of other variables.

OpenStudy (anonymous):

oh my bad different equation

OpenStudy (anonymous):

P=2l+2w when P will pass = it will be -2l+2w=-P 2w=-P+2l w=-P+2l/2